• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jmaxwell

    Log entry time 23:06:26 on April 15, 2012

    Entry number 375123

    keyword=Experiment Status

    We've had another smooth day with beautiful beam and steady target 
    polarization.  There were essentially no delays, just production all day.
    
    Production this weekend has exceeded expectations to the degree that we 
    are adding back in statistics at points that had been cut.  We're 
    approaching points which should have lower rates, so we'll up the beam 
    current as needed.
    
    If the beam current increases, another anneal may be needed tomorrow.  
    Hall C has continued to switch the half wave plate, so we haven't 
    requested it; we are continuing with positive polarization on both cups.
    
    Day Summary:
    0:00-23:00: Production